OpenStudy (anonymous):

What is the sample size of a normally distributed sample whose 100th ranked value corresponds to a standard deviate (z-score) of 0? Please round your answer to nearest integer

OpenStudy (anonymous):

Find all sides of a right triangle whose perimeter is equal to 60 cm and its area is equal to 150 square cm. A.x + y + H = 60 : perimeter , x, y and H be the two legs and the hypotenuse of the right triangle B.(1/2)xy = 150 : area C.x2 + y2 = H2: Pythagora's theorem. D.3 equations with 3 unknowns. E.(x + y)2 - 2xy = H2 : completing the square in the third equation. F.x + y = 60 - H : express x + y using the first equation and use the second equation to find xy = 300 and substitute in equation 5. G.(60 - H)2 - 600 = H2 : one equation with one unknown. H.Solve for H to find H = 25 cm. Substitute and solve for x and y to find x = 15 cm and y = 20 cm.

OpenStudy (anonymous):

What? Buddy, that has nothing to do with statistics

OpenStudy (anonymous):

srry i thought it was someone else

OpenStudy (anonymous):

I'm not sure i'm right, but standard deviate has to do with distance from the value to the mean of the sample. If the sample is normally distributed, that implies it is simmetric from the mean. So I think (since standard deviate is 0), the size of the sample must be n=200.